OpenStudy (anonymous):

Polynomial interpolation using linear systems?? Suppose we wish to find a quadratic polynomial of the form y=a_2(x^2) + a_1(x) + a_0, where a_2 a_1 and a_0 are scalars that passes through the points (-1,1),(2,4) and (p,1). Here p is a constant. Write the augmented matrix?

OpenStudy (anonymous):

Oay I don't get how you can find a_0

OpenStudy (anonymous):

thinking. joe?

OpenStudy (anonymous):

ok not thinking. looking it up

OpenStudy (anonymous):

why yes! when i see "polynomial interpolation" i think this: \[f(-1)\frac{(x-2)(x-p)}{-1-2)(-1-p)}+f(2)\frac{(x+1)(x-p)}{2+1)(2-p)}+f(p)\frac{(x+1)(x-2)}{(p+1)(p-2)}\]

OpenStudy (anonymous):

oops forgot a left parenthesis there twice >.>

OpenStudy (zarkon):

find the least squares line (quadratic)

OpenStudy (anonymous):

looks like you set up a matrix equation right? vandermonde?

OpenStudy (zarkon):

set up the standard matrix and use the normal equation \[(A^TA)^{-1}A^Tb\]

OpenStudy (anonymous):

in my book it says [1 x_1 x_1^2.... x_1^(n-1);1 x_2 x_2^2.... x_2^(n-1)] etc. and they used it for cubic polynomial

OpenStudy (anonymous):

can I use the same matrix? but the augmented matrix doesn't have the first column being fully 1's

OpenStudy (zarkon):

correct...that is the standard matrix that is needed

OpenStudy (anonymous):

no but the answer for this question doesn't have a column of 1's

OpenStudy (anonymous):

ok help me understand. because what i read had a column of 1's

OpenStudy (zarkon):

\[A=\left[\begin{matrix}1 & -1 & 1 \\ 1 & 2 & 4\\ 1 & p & p^2\end{matrix}\right]\]

OpenStudy (anonymous):

the answer is [1 -1 1 1; 4 2 1 4; p^2 p 1 1 ]?

OpenStudy (zarkon):

same thing ...I am finding \[a_0+a_1x+a_2x^2\] your matrix finds \[a_2x^2+a_1x+a_0\]

OpenStudy (zarkon):

hmm...looks like you have some extra stuff there

OpenStudy (zarkon):

an extra 1,4 and 1 which you don't need to find a quadratic polynomial

OpenStudy (anonymous):

that just comes from the: f(-1) = 1 f(2) = 4 f(p) = 1 right?

OpenStudy (anonymous):

Oh would that come from the y

OpenStudy (anonymous):

ohhhhhhhhh that make sense

OpenStudy (anonymous):

I get it now thanks for your help! but one question does it really not matter where the columns are placed

OpenStudy (anonymous):

like how the 1 column was the first and in this question it placed it in the third column

OpenStudy (zarkon):

multiply \[\left[\begin{matrix}1 & -1 & 1 \\ 1 & 2 & 4\\ 1 & p & p^2\end{matrix}\right]\left[\begin{matrix}a_0 \\a_1\\ a_2 \end{matrix}\right]\] that gives the form \[a_0+a_1x+a_2x^2\]

OpenStudy (zarkon):

now compute ... \[(A^TA)^{-1}A^Tb\] b being the vector consisting of the y-values

OpenStudy (anonymous):

Okay my only question was why my answer key put it in a_2x^2 + a_1x + a_0

OpenStudy (zarkon):

@joe yes...but his matrix had it in twice which is not correct.

OpenStudy (zarkon):

that way is fine..it is really the same form

OpenStudy (anonymous):

No I understand what joe is saying, it make sense, okay thanks for your help everyone

OpenStudy (zarkon):

I get \[\left(\frac{2}{p-2}+2\right)x^2+\frac{p-1}{p-2}x-\frac{1}{p-2}\]

OpenStudy (anonymous):

my other question is why is it infinitely solutions when p=-1

OpenStudy (anonymous):

the matrix when reduced is [1 0 0 -1/(p-2); 0 1 0 (p-1)/(p-2); 0 0 1 2(p-1)/(p-2)]

OpenStudy (anonymous):

if be was -1, then the first and third row of the matrix would be the same, and it would be singular.

OpenStudy (zarkon):

same problem with 2

OpenStudy (anonymous):

ohh I that what it means by infinitely many I thought it was when 1 =1 or something

OpenStudy (anonymous):

so if two rows are the same then the solutions are infinite

OpenStudy (zarkon):

then the matrix will not have full rank and therefore \[A^TA\] will not be invertable

OpenStudy (anonymous):

okay make sense, thanks again! I'm actually liking this stuff b/c it make sense now haha

OpenStudy (zarkon):

cool

OpenStudy (anonymous):

btw, that formula i put at the top also gives that solution you got by calculating: \[(A^TA)^{-1}A^Tb\] just thought i would point that out lol

OpenStudy (anonymous):

after you simplify it and whatnot.

OpenStudy (zarkon):

is that the Lagrange polynomial

OpenStudy (anonymous):

i think it is..i caught a random glimpse of it in a book once. Let me see if i can find it....

OpenStudy (zarkon):

I'm pretty sure it is...I'd be a little surprised if it always gave the least squares solution

OpenStudy (anonymous):

yeah, its called Lagrange Interpolation in this book.

OpenStudy (zarkon):

if we had more than 3 points then it would be different

OpenStudy (anonymous):

i remember reading this section and not understanding a lot of it >.< it just came to mind because of the word "interpolation" lol >.<